OpenStudy (anonymous):

Divide. Write your answer in simplest form. 5/8 divided by 3/10

OpenStudy (anonymous):

5/8 * 10/3 = 50/21 then you divide

OpenStudy (anonymous):

it needs to be a fraction though

OpenStudy (anonymous):

Ok so multiply 8/5 times 3/10

OpenStudy (anonymous):

so 24/50

OpenStudy (anonymous):

yes and simplify

OpenStudy (anonymous):

I got 25/12 or 2 1/12.

OpenStudy (anonymous):

I got 12/25

OpenStudy (anonymous):

it says its wrong

OpenStudy (anonymous):

Hmmm... Here I am going to put it into a generator site and check who is right

OpenStudy (anonymous):

Im wrong ts 15/80

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

@NatPanda http://www.wolframalpha.com/input/?i=5%2F8+divided+by+3%2F10 I would trust this, but I will keep checking for you

OpenStudy (anonymous):

Because I learned it as keep change flip or KCF so you keep the first fraction then change the division sign and flip the last fraction so it would be: 5/8 x 3/10= 15/80= 3/16 the solution is 3/16

OpenStudy (anonymous):

I don't think that is right still @Thomas2345. Idk maybe I'm off the peg lol